Cliff is 6 years older than his sister Claire. In 2 years cliff will be twice as old.
X = Cliff present Age
Y = Claire's present Age
Question: what is their current age?

Answer by stanbon(75887)

Cliff is 6 years older than his sister Claire.
In 2 years cliff will be twice as old.
---
Equations:
X = Cliff present Age
Y = Claire's present Age
x = y+6
(x+2) = 2(y+2)
----
substitute for "x" and solve for "y":
(y+6+2) = 2y+4
y + 8 = 2y+4
y = 4 (Claire's present age)
----
Solve for "x":
x = y + 6
x = 4 + 6 = 10 (Cliff's present age)
=======================================
Cheers,
Stan H.
==========
-

Question: what is their current age?